If an electron is accelerated from rest through a potential difference of 12.0 kV, what is its resulting speed? (e = 1.60 × 10-19 C, k = 1/4πε0 = 8.99 × 109 N ∙ m2/C2, mel = 9.11 x 10-31 kg). (Give your answer to the nearest km/s).
Answer #1

given

   potential differnce V = 12.0 kV = 12000 V.

charge of electron e = -1.6*10^-19 C, k = 1/4πε0 = 8.99 × 109 N ∙ m2/C2, mel = 9.11 x 10-31 kg

we know that the potential difference is = W/e = change in k.e / e


electron is accelerated from rest so initila velocity is 0 m/s, final velocity is V

change in k.e = 1/2 m v^2

       12000 = 0.5*9.11*10^-31*v^2 / 1.6*10^-19

solving for v = 64924172.60m/s = 64924.17260 km/s

speed of electron is 64924.1726 km/s